# PolyOCP.jl– A Julia Package for Stochastic OCPs and MPC

### v2

@@ Abstract
The consideration of stochastic uncertainty in optimal and predictive control is a well-explored topic. Recently Polynomial Chaos Expansions (PCE) have seen a lot of considerations for problems involving stochastically uncertain system parameters and also for problems with additive stochastic i.i.d. disturbances. While there exist a number of open-source PCE toolboxes, tailored open-source codes for the solution of OCPs involving additive stochastic i.i.d. disturbances in julia are not available. Hence, this paper introduces the toolbox PolyOCP.jl which enables to efficiently solve stochastic OCPs for a large class of disturbance distributions. We explain the main mathematical concepts between the PCE transcription of stochastic OCPs and how they are provided in the toolbox. We draw upon two examples to illustrate the functionalities of PolyOCP.jl.
@@ I Introduction
The consideration of stochastic uncertainty in optimal and predictive control is a well-explored topic. Landmark results range from Witsenhausen’s counterexample, which shows that in non-Gaussian time-varying linear-quadratic output-feedback settings the optimal policy might be non-linear [1], through stochastic extensions of the LQR [2, 3] to recent investigations of stochastic MPC [4, 5, 6, 7, 8] and the scenario approach towards handling uncertainty [9]. Indeed, in settings with additive Gaussian uncertainty one may take the classic LQG route [10]. When it comes to non-Gaussian disturbances (of finite expectation and variance) and LTI systems, the propagation of first and second moments is structurally identical to the Gaussian setting. For a detailed comparison of different approaches to uncertainty propagation for dynamic systems, we refer to [11]. However, considering only the first two moments may cause a significant loss of distributional information, e.g., for formulations with non-quadratic cost functions [12] and for chance constraints reformulations involving higher-order moments [13]. Hence, methods that can work with complete distributional information are of interest.
@@
Indeed there exist a number of toolboxes which facilitate PCE, see Table I, and also [19], for an overview. PoCET [20] addresses Galerkin projection for PCE and parametric uncertainty in dynamic systems. PolyChaos.jl provides a framework for efficient construction of PCE bases using quadrature rules and tensorized computation in julia. For instance, PolyChaos.jl includes an example on how to solve stochastic OCP with parametrically uncertain systems. However, the packages focus on generating the PCE of parameters and do not provide functions specifically for stochastic control. Interestingly, none of these toolboxes is explicitly conceived to handle additive disturbances acting on a dynamic system. Put differently, none of the toolboxes mentioned in Table I is tailored to simplify the solution of stochastic Optimal Control Problems (OCP) with additive uncertainty. One reason is that additive i.i.d. disturbance processes acting on dynamic systems require to expand the dimension of the PCE basis as the horizon grows, see, e.g., [21] for a PCE-based analysis of LQ optimal control.
@@
The remainder of the paper is structured as follows: Section II provides background on the PCE transcription of stochastic OCPs, while Section III introduces core functionalities of the PolyOCP.jl. Section IV presents selected examples. The paper ends with conclusions and outlook in Section V.
@@ II Problem Formulation
Next, we introduce the problem formulation and provide some background on PCE and stochastic OCPs.
@@ II-A Polynomial chaos expansions
The key idea of PCE is that any $\mathcal{L}^{2}$ random variable can be expressed in a suitable orthogonal polynomial basis. Consider a univariate orthogonal polynomial basis $\{\phi^{j}(\xi)\}_{j=0}^{\infty}$ that spans the space $\mathcal{L}^{2}(\Xi,\mathcal{F},\mu;\mathbb{R})$, where $\phi^{j}(\xi)$ is a polynomial function of degree $j$, $\Xi$ the sample space of $\xi$, $\mathcal{F}$ a $\sigma$-algebra, and $\mu$ the considered probability measure. Note that $\xi\in\mathcal{L}^{2}(\Xi,\mathcal{F},\mu;\mathbb{R})$ is the stochastic germ of polynomial functions. The orthogonality of $\{\phi^{j}(\xi)\}_{j=0}^{\infty}$ ensures that
$\langle\phi^{i}(\xi),\phi^{j}(\xi)\rangle{=}\int_{\Xi}\phi^{i}(\xi)\phi^{j}(\xi)\mathop{}\!\mathrm{d}\mu(\xi){=}\delta^{ij}\langle\phi^{j}(\xi),\phi^{j}(\xi)\rangle$ (1)
holds, where $\delta^{ij}$ denotes the Kronecker delta. The first polynomial $\phi^{0}(\xi)$ is of degree 0 and thus always chosen to be $\phi^{0}(\xi)=1$. Hence, the orthogonality (1) gives that for all other basis dimensions $j>0$, we have $\mathbb{E}[\phi^{j}(\xi)]=\int_{\Xi}\phi^{j}(\xi)\mathop{}\!\mathrm{d}\mu(\xi)=\langle\phi^{j}(\xi),\phi^{0}(\xi)\rangle=0$.
@@
where $\textsf{z}^{j}\in\mathbb{R}$ is referred to as the $j$-th PCE coefficient. In the PCE expression, $\xi:\Omega\to\Xi$ is viewed as a function of the outcome $\omega$ and hence $\phi^{j}(\xi(\omega))=\phi^{j}\circ\xi(\omega)$. For the sake of readability, we omit the arguments $\xi(\omega)$, $\omega$ and use the shorthand $\mathcal{L}^{2}(\mathbb{R})\coloneqq\mathcal{L}^{2}(\Omega,\mathcal{F},\mu;\mathbb{R})$ whenever there is no ambiguity. The first two moments of $Z$ thus can be efficiently computed from its PCE coefficients as
$\mathbb{E}[Z]=\textsf{z}^{0},\quad\mathbb{V}[Z]=\sum_{j=1}^{\infty}\textsf{z}^{2}\langle\phi^{j},\phi^{j}\rangle.$ (2)
@@
where $L\in\mathbb{N}^{\infty}\coloneqq\mathbb{N}^{+}\cup\{\infty\}$ is the PCE dimension. For $L\to\infty$, the truncation error satisfies $\lim_{L\to\infty}\|\Delta Z(L)||=0$ [32, 33].
@@
The correspondence between Wiener-Askey polynomial families and their underlying continuous stochastic germs is summarized in Table II, which serves as a guideline for selecting the appropriate polynomial basis for random variables with canonical distributions [34]. The Dirac measure, representing a deterministic variable with polynomial basis $\phi=1$, is also included to foster modelling freedom. Furthermore, Table II lists the first two nontrivial polynomials (beyond $\phi^{0}=1$) for each distribution type, along with their normalized counterparts $\psi^{j}$ satisfying $\langle\psi^{j},\psi^{j}\rangle=1$, obtained via
@@
Given an $\mathcal{L}^{2}$ random variable with known distribution, the key to constructing an exact finite-dimensional PCE is the appropriate choice of basis functions. For some widely used distributions, the appropriate choice of polynomial bases is summarized in Table II [34]. Additionally, a generic (non-orthonormal but orthogonal) basis choice for any random variable $Z\in\mathcal{L}^{2}(\mathbb{R})$ is $\phi^{0}=1$ and $\phi^{1}=Z-\mathbb{E}[Z]$, which implies the exact and finite PCE $\textsf{z}^{0}=\mathbb{E}[Z]$ and $\textsf{z}^{1}=1$. $\square$
@@
Distribution of $Z$ Distribution of $\xi$ basis Orthogonal Basis $\phi^{j}$ Orthonormal Basis $\psi^{j}$ Juila type in PolyOCP.jl
Dirac: $\delta_{x}(c)$ $\delta_{x}(0)$ Constant $\phi^{0}=1$ $\psi^{0}=\phi^{0}$ DiracMeasureParametric
Gaussian: $\mathcal{N}(\mu,\sigma^{2})$ $\mathcal{N}(0,1)$ Hermite $He_{j}(\xi)$ $\phi^{1}(\xi)=\xi$ $\psi^{1}=\phi^{1}$ GaussMeasureParametric
$\phi^{2}(\xi)=\xi^{2}-1$ $\psi^{2}(\xi)=\phi^{2}/\sqrt{2}$
Uniform: $\mathcal{U}(0,1)$ $\mathcal{U}(0,1)$ Legendre $P_{j}(2\xi-1)$ $\phi^{1}(\xi)=\xi-\tfrac{1}{2}$ $\psi^{1}=2\sqrt{3}\phi^{1}$ UniformMeasureParametric
$\phi^{2}(\xi)=\xi^{2}-\xi+\tfrac{1}{6}$ $\psi^{2}=6\sqrt{5}\phi^{2}$
Beta: $\mathcal{B}(\alpha{=}2,\beta{=}2)$ $\mathcal{B}(\alpha,\beta)$ Jacobi $J_{j}^{(\alpha-1,\beta-1)}(2\xi-1)$ $\phi^{1}(\xi)=\xi-\tfrac{1}{2}$ $\psi^{1}=2\sqrt{5}\phi^{1}$ BetaMeasureParametric
$\phi^{2}(\xi)=\xi^{2}-\xi+\tfrac{1}{5}$ $\psi^{2}=5\sqrt{14}\phi^{2}$
Gamma: $\Gamma(k{=}1,\theta{=}1)$ $\Gamma(k,\theta)$ Laguerre $L_{j}(\xi)$ $\phi^{1}(\xi)=\xi-1$ $\psi^{1}=\phi^{1}$ GammaMeasureParametric
$\phi^{2}(\xi)=\xi^{2}-4\xi+2$ $\psi^{2}=\phi^{2}/2$
@@
$\begin{split}\{\phi(\xi)\}_{j=0}^{\infty}&=\{1\}\cup\Phi^{1}(\xi)\cup\Phi^{2}(\xi)\cup\cdots\cup\Phi^{\infty}(\xi),\\ \text{with}\quad\Phi^{p}(\xi)&\coloneqq\bigg\{\prod_{i=0}^{{n_{z}}}\phi_{i}^{j_{i}}(\xi_{i})\,\Big|\,j_{i}\in\mathbb{I}_{[0,j]},\,\sum_{i=1}^{{n_{z}}}j_{i}=p\bigg\},\end{split}$ (3)
@@
$\Phi^{1}(\xi)\coloneqq$ $\{\phi_{1}^{1}(\xi_{1}),\phi_{1}^{1}(\xi_{2}),\cdots,\phi_{{n_{z}}}^{1}(\xi_{{n_{z}}})\},$
$\Phi^{2}(\xi)\coloneqq$ $\Big\{\phi_{1}^{2}(\xi_{1}),\cdots,\phi_{{n_{z}}}^{2}(\xi_{{n_{z}}}),$
$\quad,\phi_{1}^{1}(\xi_{1})\cdot\phi_{2}^{1}(\xi_{2}),\cdots,\phi_{{n_{z}}}^{1}(\xi_{{n_{z}}-1})\cdot\phi_{{n_{z}}}^{1}(\xi_{{n_{z}}})\Big\}.$
@@
where the stochastic germs $\xi_{1}\sim\mathcal{U}(0,1)$, $\xi_{2}\sim\mathcal{N}(0,1)$ are independent. By constructing a multivariate basis $\{1,\phi_{1}^{1}(\xi_{1}),\phi_{2}^{1}(\xi_{2})\}$, the PCE of Z reads
@@
where $\mathrm{Chol}(\Sigma)\in\mathbb{R}^{{n_{z}}\times{n_{z}}}$ denotes a lower-triangular matrix obtained by Cholesky decomposition of $\Sigma$ [27]. The basis functions are chosen as $\phi_{i}^{1}(\xi_{i})=\xi_{i}$ with independent stochastic germs $\xi_{i}\sim\mathcal{N}(0,1)$ for all $i\in\mathbb{I}_{[1,{n_{z}}]}$. $\square$
Note that the multivariate bases in the above examples preserve orthogonality because the stochastic germs are independent random variables. The multivariate Gaussian distribution in Example 2 allows addressing distributionally robust uncertainty propagation within the PCE framework [27].
@@ II-B Uncertainty propagation
Importantly, (5) indicates that $Y$ can be exactly represented in the basis of $Z$ and thus no new terms are introduced. This is in contrast to nonlinear maps as demonstrated in Example 3.
@@ II-C Stochastic optimal control
where $Q$, $Q_{N}\succeq 0$, $R\succ 0$, $\|X(k)\|_{Q}^{2}\coloneqq X^{\top}(k)QX(k)\in\mathcal{L}^{2}(\mathbb{R})$, and $(z,Z,N^{\prime})\in\{(x,X,N),\,(u,U,N-1)\}$.
@@
The initial condition $X_{\text{ini}}$ and all i.i.d. disturbances $W_{k}$, $k\in\mathbb{I}_{[0,N-1]}$ admit exact PCEs with $L_{\text{ini}}$ terms and $L_{w}$ terms of orthonormal bases, respectively. That is, $X_{\text{ini}}=\textstyle{\sum_{i=0}^{L_{\text{ini}}-1}}\textsf{x}_{\text{ini}}^{i}\psi_{\text{ini}}^{i}(\xi_{\text{ini}})$ and $W_{k}=\textstyle{\sum_{n=0}^{L_{w}-1}}\textsf{w}_{k}^{n}\psi_{w}^{n}(\xi_{k})$ for $k\in\mathbb{I}_{[0,N-1]}$, where $\xi_{k}$ are i.i.d. stochastic germs.
In the above assumption, each $\xi_{k}$, $k\in\mathbb{I}_{[0,N-1]}$ corresponds to the disturbance $W(k)$ at time step $k$. In the PCEs of i.i.d. disturbances, the identical distribution is expressed via the shared algebraic structure of the basis functions $\psi_{w}$, while the stochastic independence is modeled by the use of different germs $\xi_{k}$. Similar to the multivariate basis illustrated in Example 1, we construct a joint orthonormal basis $\Psi=\{\psi^{j}(\xi)\}_{j=0}^{L-1}$ with $\xi=\begin{bmatrix}\xi_{\text{ini}}&\xi_{1}&\cdots\xi_{N-1}\end{bmatrix}^{\top}$ for $X_{\text{ini}}$ and $W(k)$, $k\in\mathbb{I}_{[0,N-1]}$ as
$\Psi\coloneqq\{\psi_{\text{ini}}^{j}(\xi_{\text{ini}})\}_{i=0}^{L_{\text{ini}}-1}\bigcup\left(\bigcup_{k=0}^{N-1}\{\psi_{w}^{n}(\xi_{k})\}_{n=0}^{L_{w}-1}\right)$ (8)
@@
with a total of $L=L_{\text{ini}}+N(L_{w}-1)$ terms, i.e., it grows linearly with the horizon $N$. Under suitable technical conditions, it can be shown that the optimal solution of (7), i.e. $\{X^{\star}(k)\}_{k=0}^{N}$ and $\{U^{\star}(k)\}_{k=0}^{N-1}$, admits exact PCEs in the joint basis $\Psi$ from (8), see [36] for details.
By replacing all the random variables in system (6) with their PCEs using the joint basis $\Psi$, e.g. $X(k)=\sum_{j=0}^{L-1}\textsf{x}^{j}(k)\psi^{j}$, the resulting system for PCE coefficients satisfies
@@
Using Cantelli’s inequality, a conservative component-wise approximation of the chance constraint (7c) is given by
$\mathbb{E}[Z]\pm\gamma(\varepsilon_{z})\mathbb{V}[Z]\in[z_{\min},z_{\max}]\quad\text{with}\quad\gamma(\varepsilon_{z})=\sqrt{\frac{2-\varepsilon_{z}}{\varepsilon_{z}}},$
@@
PolyOCP.jl Stochastic OCP (7) Stochastic OCP (10) PCE transcription PolyChaos.jl PolyOCP.jl JuMP.jl PCE solution $\textsf{u}^{j,\star}(k),\,\textsf{x}^{j,\star}(k)$
Figure 1: Flow chart of solving stochastic OCPs using PolyOCP.jl
@@
Getting PCE representations of random variables serves as the first step in solving stochastic OCPs within the PCE framework, i.e.,  (10). Building upon PolyChaos.jl, PolyOCP.jl defines the canonical parametric measures and their corresponding orthonormal polynomial functions, as summarized in Table II. Note that the use of orthonormal bases in PolyOCP.jl simplifies the formulation of the objective (10a) and the chance constraint (10c), and also improves numerical stability. Given a random vector $Z\in\mathcal{L}^{2}({\mathbb{R}^{{n_{z}}})}$ whose components are defined as types listed in Table II, the function genPCE directly returns the sparse PCE representation of $Z$, including the multivariate orthonormal basis (MultiOrthonoPoly) and the associated coefficients. Moreover, PolyOCP.jl supports user-defined PCE of $Z$ with multivariate bases up to arbitrary degrees and specified PCE coefficients.

### v3

@@ Abstract
The consideration of stochastic uncertainty in optimal and predictive control is a well-explored topic. Recently Polynomial Chaos Expansions (PCE) have received considerable attention for problems involving stochastically uncertain system parameters and also for problems with additive stochastic i.i.d. disturbances. While there exist a number of open-source PCE toolboxes, tailored open-source codes for the solution of OCPs involving additive stochastic i.i.d. disturbances in julia are not available. Hence, this paper introduces the toolbox PolyOCP.jl which enables to efficiently solve stochastic OCPs for linear systems subject to a large class of disturbance distributions. We explain the main mathematical concepts between the PCE transcription of stochastic OCPs and how they are provided in the toolbox. We draw upon two examples to illustrate the functionalities of PolyOCP.jl.
@@ I Introduction
The consideration of stochastic uncertainty in optimal and predictive control is a well-explored topic. Landmark results range from Witsenhausen’s counterexample, which shows that in non-Gaussian time-varying linear-quadratic output-feedback settings the optimal policy might be non-linear [1], through stochastic extensions of the LQR [2, 3] to recent investigations of stochastic MPC [4, 5, 6, 7, 8] and the scenario approach for handling uncertainty [9]. Indeed, in settings with additive Gaussian uncertainty one may take the classic LQG route [10]. When it comes to non-Gaussian disturbances (of finite expectation and variance) and LTI systems, the propagation of first and second moments is structurally identical to the Gaussian setting. For a detailed comparison of different approaches to uncertainty propagation for dynamic systems, we refer to [11]. However, considering only the first two moments may cause a significant loss of distributional information, e.g., for formulations with non-quadratic cost functions [12] and for chance constraints reformulations involving higher-order moments [13]. Hence, methods that can work with complete distributional information are of interest.
@@
Indeed there exist a number of toolboxes which facilitate PCE, see Table I, and also [19], for an overview. PoCET [20] addresses Galerkin projection for PCE and parametric uncertainty in dynamic systems. PolyChaos.jl [21] provides a framework for efficient construction of PCE bases using quadrature rules and tensorized computation in julia. For instance, PolyChaos.jl includes an example on how to solve stochastic OCP with parametrically uncertain systems. However, the packages focus on generating the PCE of parameters and do not provide functions specifically for stochastic control. Interestingly, none of these toolboxes is explicitly conceived to handle additive disturbances acting on a dynamic system. Put differently, none of the toolboxes mentioned in Table I is tailored to simplify the solution of stochastic Optimal Control Problems (OCP) with additive uncertainty. One reason is that additive i.i.d. disturbance processes acting on dynamic systems require to expand the dimension of the PCE basis as the horizon grows, see, e.g., [22] for a PCE-based analysis of LQ optimal control.
@@
The remainder of the paper is structured as follows: Section II provides background on the PCE transcription of stochastic OCPs, while Section III introduces core functionalities of the PolyOCP.jl. Section IV presents selected examples. The paper concludes with Section V.
@@ II Problem Formulation
Next, we provide some background on PCE and stochastic OCPs and introduce the problem formulation.
@@ II-A Polynomial chaos expansions
The key idea of PCE is that any $\mathcal{L}^{2}$ random variable can be expressed in a suitable orthogonal polynomial basis. Consider a univariate orthogonal polynomial basis $\{\phi^{j}(\xi)\}_{j=0}^{\infty}$ that spans the space $\mathcal{L}^{2}(\Omega,\mathcal{F},\mu;\mathbb{R})$, where $\phi^{j}(\xi)$ is a polynomial of degree jj [^1] In general, one may also consider different indexing schemes where $j$ does not correspond to the degree of $\phi^{j}$., $\xi\in\mathcal{L}^{2}(\Omega,\mathcal{F},\mu;\Xi)$ is the argument of polynomials. Here, $\Omega$ is the sample space, $\mathcal{F}$ a $\sigma$-algebra, $\mu$ the considered probability measure. Note that $\xi:\Omega\to\Xi$ is viewed as a function of the outcome $\omega$ and hence $\phi^{j}(\xi(\omega))=\phi^{j}\circ\xi(\omega)$. The orthogonality of $\{\phi^{j}(\xi)\}_{j=0}^{\infty}$ ensures that
$\hskip-2.84526pt\langle\phi^{i}(\xi),\phi^{j}(\xi)\rangle{=}\int_{\Omega}\hskip-2.84526pt\phi^{i}(\xi)\phi^{j}(\xi)\mathop{}\!\mathrm{d}\mu{=}\delta^{ij}\langle\phi^{j}(\xi),\phi^{j}(\xi)\rangle$ (1)
holds, where $\delta^{ij}$ denotes the Kronecker delta. The first polynomial $\phi^{0}(\xi)$ is of degree 0 and thus always chosen to be $\phi^{0}(\xi)=1$. Hence, the orthogonality (1) gives that for all other basis dimensions $j>0$, we have $\mathbb{E}[\phi^{j}(\xi)]=\int_{\Omega}\phi^{j}(\xi)\mathop{}\!\mathrm{d}\mu=\langle\phi^{j}(\xi),\phi^{0}(\xi)\rangle=0$.
@@
where $\textsf{z}^{j}\in\mathbb{R}$ is referred to as the $j$-th PCE coefficient. For the sake of readability, we omit the arguments $\xi(\omega)$, $\omega$ and use the shorthand $\mathcal{L}^{2}(\mathbb{R})\coloneqq\mathcal{L}^{2}(\Omega,\mathcal{F},\mu;\mathbb{R})$ whenever there is no ambiguity. The first two moments of $Z$ thus can be efficiently computed from its PCE coefficients as
$\mathbb{E}[Z]=\textsf{z}^{0},\quad\mathbb{V}[Z]=\sum\nolimits_{j=1}^{\infty}(\textsf{z}^{j})^{2}\langle\phi^{j},\phi^{j}\rangle.$ (2)
@@
where $L\in\mathbb{N}^{\infty}\coloneqq\mathbb{N}^{+}\cup\{\infty\}$ is the PCE dimension. For $L\to\infty$, the truncation error satisfies $\lim_{L\to\infty}\|\Delta Z(L)\|=0$ [32, 33].
@@
The correspondence between Wiener-Askey polynomial families and their underlying continuous stochastic arguments is summarized in Table II, which serves as a guideline for selecting the appropriate polynomial basis for random variables with canonical distributions [34]. The Dirac measure, representing a deterministic variable with basis $\phi=1$, is also included to foster modelling freedom. Furthermore, Table II lists the first two nontrivial polynomials (beyond $\phi^{0}=1$) for each distribution type, along with their normalized counterparts $\psi^{j}$ satisfying $\langle\psi^{j},\psi^{j}\rangle=1$, obtained via
@@
Given an $\mathcal{L}^{2}$ random variable with known distribution, the key to constructing an exact finite-dimensional PCE is the appropriate choice of basis functions. For some widely used distributions, the appropriate choice of polynomial bases and the corresponding affine PCE expansions are summarized in Table II [34]. Additionally, a generic (non-orthonormal but orthogonal) basis choice for any random variable $Z\in\mathcal{L}^{2}(\mathbb{R})$ is $\phi^{0}=1$ and $\phi^{1}=Z-\mathbb{E}[Z]$, which implies the exact and finite PCE $\textsf{z}^{0}=\mathbb{E}[Z]$ and $\textsf{z}^{1}=1$. $\square$
@@
Distribution of $Z$ Distribution of $\xi$ Basis Orthogonal Basis $\phi^{j}$ Orthonormal Basis $\psi^{j}$ Affine PCE
Dirac: $\delta_{x}(c)$ $\delta_{x}(0)$ Constant $\phi^{0}=1$ $\psi^{0}=\phi^{0}$ $Z=c\cdot\phi^{0}=c\cdot\psi^{0}$
Gaussian: $\mathcal{N}(\mu,\sigma^{2})$ $\mathcal{N}(0,1)$ Hermite $\phi^{1}(\xi)=\xi$ $\psi^{1}=\phi^{1}$ $Z=\mu\cdot\phi^{0}+\sigma\cdot\phi^{1}$
$\phi^{2}(\xi)=\xi^{2}-1$ $\psi^{2}(\xi)=\phi^{2}/\sqrt{2}$ $\phantom{Z}=\mu\cdot\psi^{0}+\sigma\cdot\psi^{1}$
Uniform: $\mathcal{U}(a,b)$ $\mathcal{U}(0,1)$ Legendre $\phi^{1}(\xi)=\xi-\tfrac{1}{2}$ $\psi^{1}=2\sqrt{3}\phi^{1}$ $Z=\frac{a+b}{2}\cdot\phi^{0}+(b{-}a)\cdot\phi^{1}$
$\phi^{2}(\xi)=\xi^{2}-\xi+\tfrac{1}{6}$ $\psi^{2}=6\sqrt{5}\phi^{2}$ $\phantom{Z}=\frac{a+b}{2}\cdot\psi^{0}+\frac{b{-}a}{2\sqrt{3}}\cdot\psi^{1}$
Beta: $\mathcal{B}(\alpha{=}2,\beta{=}2)$ $\mathcal{B}(\alpha,\beta)$ Jacobi $\phi^{1}(\xi)=\xi-\tfrac{1}{2}$ $\psi^{1}=2\sqrt{5}\phi^{1}$ $Z=\frac{1}{2}\cdot\phi^{0}+1\cdot\phi^{1}$
$\phi^{2}(\xi)=\xi^{2}-\xi+\tfrac{1}{5}$ $\psi^{2}=5\sqrt{14}\phi^{2}$ $\phantom{Z}=\frac{1}{2}\cdot\psi^{0}+\frac{1}{2\sqrt{5}}\cdot\psi^{1}$
Gamma: $\Gamma(\alpha{=}1,\beta{=}1)$ $\Gamma(\alpha,\beta)$ Laguerre $\phi^{1}(\xi)=\xi-1$ $\psi^{1}=\phi^{1}$ $Z=1\cdot\phi^{0}+1\cdot\phi^{1}$
$\phi^{2}(\xi)=\xi^{2}-4\xi+2$ $\psi^{2}=\phi^{2}/2$ $\phantom{Z}=1\cdot\psi^{0}+1\cdot\psi^{1}$
@@
$\begin{split}\{\phi(\xi)\}_{j=0}^{\infty}&=\{1\}\cup\Phi^{1}(\xi)\cup\Phi^{2}(\xi)\cup\cdots\cup\Phi^{\infty}(\xi),\\ \text{with}\quad\Phi^{p}(\xi)&\coloneqq\bigg\{\prod_{i=0}^{{n_{z}}}\phi_{i}^{j_{i}}(\xi_{i})\,\Big|\,j_{i}\in\mathbb{I}_{[0,p]},\,\sum_{i=1}^{{n_{z}}}j_{i}=p\bigg\},\end{split}$ (3)
@@
$\Phi^{1}(\xi)$ $\coloneqq\{\phi_{1}^{1}(\xi_{1}),\phi_{1}^{1}(\xi_{2}),\ldots,\phi_{{n_{z}}}^{1}(\xi_{{n_{z}}})\},$
$\Phi^{2}(\xi)$ $\coloneqq\Big\{\phi_{1}^{2}(\xi_{1}),\ldots,\phi_{{n_{z}}}^{2}(\xi_{{n_{z}}}),$
$\hskip 17.07164pt\phi_{1}^{1}(\xi_{1})\cdot\phi_{2}^{1}(\xi_{2}),\ldots,\phi_{{n_{z}}-1}^{1}(\xi_{{n_{z}}-1})\cdot\phi_{{n_{z}}}^{1}(\xi_{{n_{z}}})\Big\}.$
@@
where the stochastic arguments $\xi_{1}\sim\mathcal{U}(0,1)$, $\xi_{2}\sim\mathcal{N}(0,1)$ are independent. By constructing a multivariate basis $\{1,\phi_{1}^{1}(\xi_{1}),\phi_{2}^{1}(\xi_{2})\}$, the PCE of Z reads
@@
where $\mathrm{Chol}(\Sigma)\in\mathbb{R}^{{n_{z}}\times{n_{z}}}$ denotes a lower-triangular matrix obtained by Cholesky decomposition of $\Sigma$ [27]. The basis functions are chosen as $\phi_{i}^{1}(\xi_{i})=\xi_{i}$ with independent stochastic arguments $\xi_{i}\sim\mathcal{N}(0,1)$ for all $i\in\mathbb{I}_{[1,{n_{z}}]}$. $\square$
Note that the multivariate bases in the above examples preserve orthogonality because the stochastic arguments are independent random variables. The multivariate Gaussian distribution in Example 2 allows addressing distributionally robust uncertainty propagation within the PCE framework [27].
@@ II-B Uncertainty propagation
Importantly, (5) indicates that $Y$ can be exactly represented in the basis of $Z$ and thus no new terms are introduced. This is in contrast to nonlinear maps, cf. Example 3.
@@ II-C Stochastic optimal control
where $Q$, $Q_{N}\succeq 0$, $R\succ 0$, $\|X(k)\|_{Q}^{2}\coloneqq X^{\top}(k)QX(k)\in\mathcal{L}^{2}(\mathbb{R})$, and $(z,Z,N^{\prime})\in\{(x,X,N),\,(u,U,N-1)\}$. Note that we have $k\in\mathbb{I}_{[0,N]}$ for the state constraints, while $k\in\mathbb{I}_{[0,N-1]}$ for the input constraints, reflecting that the predicted state trajectory is one step longer than the control trajectory.
@@
The initial condition $X_{\text{ini}}$ and all i.i.d. disturbances $W_{k}$, $k\in\mathbb{I}_{[0,N-1]}$ admit exact PCEs with $L_{\text{ini}}$ terms and $L_{w}$ terms of orthonormal bases, respectively. That is, $X_{\text{ini}}=\textstyle{\sum_{i=0}^{L_{\text{ini}}-1}}\textsf{x}_{\text{ini}}^{i}\psi_{\text{ini}}^{i}(\xi_{\text{ini}})$ and $W_{k}=\textstyle{\sum_{n=0}^{L_{w}-1}}\textsf{w}_{k}^{n}\psi_{w}^{n}(\xi_{k})$ for $k\in\mathbb{I}_{[0,N-1]}$, where $\xi_{k}$ are i.i.d. stochastic arguments.
[…]
$\Psi\coloneqq\{\psi_{\text{ini}}^{i}(\xi_{\text{ini}})\}_{i=0}^{L_{\text{ini}}-1}\bigcup\left(\bigcup_{k=0}^{N-1}\{\psi_{w}^{n}(\xi_{k})\}_{n=0}^{L_{w}-1}\right)$ (8)
@@
with a total of $L=L_{\text{ini}}+N(L_{w}-1)$ terms, i.e., it grows linearly with the horizon $N$. Sufficient conditions ensuring that the optimal solution of (7), i.e. $\{X^{\star}(k)\}_{k=0}^{N}$ and $\{U^{\star}(k)\}_{k=0}^{N-1}$, admits exact PCEs in the joint basis $\Psi$ from (8) have been analyzed in [36].
By replacing all the random variables in (6) with their PCEs using the joint basis $\Psi$, e.g. $X(k)=\sum_{j=0}^{L-1}\textsf{x}^{j}(k)\psi^{j}$, the resulting system for PCE coefficients satisfies
@@
Using Cantelli’s inequality, a conservative component-wise approximation of the chance constraint (7c) is
$\mathbb{E}[Z]\pm\gamma(\varepsilon_{z})\sqrt{\mathbb{V}[Z]}\in[z_{\min},z_{\max}]\;\text{with}\;\gamma(\varepsilon_{z})=\sqrt{\frac{2-\varepsilon_{z}}{\varepsilon_{z}}},$
@@
By solving the reformulated stochastic OCP (10), one obtains the PCE expansion of the optimal input $U^{\star}(k;\omega)=\sum_{j=0}^{L-1}\textsf{u}^{j}(k)\psi^{j}(\xi(\omega))$. Note that the realizations of the basis functions $\psi(\xi(\omega))$, $j\in\mathbb{I}_{[0,L-1]}$ are determined by the disturbance realizations $W(k;\omega)$, $k\in\mathbb{I}_{[0,N-1]}$. Hence, the optimal input $U^{\star}$ obtained by solving the open-loop OCP (10) is in fact a feedback policy with respect to future disturbance realizations. Moreover, the optimal input is not necessarily an affine feedback as long as it lives in the space spanned by the PCE basis $\Psi=\{\psi^{j}\}_{j=0}^{L-1}$.
@@
PolyOCP.jl Stochastic OCP (7) Stochastic OCP in PCE (10) PCE transcription PolyChaos.jl PolyOCP.jl JuMP.jl PCE solution $\textsf{u}^{j,\star}(k),\,\textsf{x}^{j,\star}(k)$
Figure 1: Solving stochastic OCPs using PolyOCP.jl
@@
Getting PCE representations of random variables serves as the first step in solving stochastic OCPs within the PCE framework, i.e. (10). Building upon PolyChaos.jl, PolyOCP.jl defines the canonical parametric measures and their corresponding orthonormal polynomial functions, as summarized in Table II. Note that the use of orthonormal bases in PolyOCP.jl simplifies the formulation of the objective (10a) and the chance constraint (10c), and also improves numerical stability. Given a random vector $Z\in\mathcal{L}^{2}({\mathbb{R}^{{n_{z}}})}$, whose components follow the distributions listed in Table II, the function genPCE directly returns the sparse PCE representation of $Z$, including the multivariate orthonormal basis (MultiOrthonoPoly) and the associated coefficients. Moreover, PolyOCP.jl supports user-defined PCE of $Z$ with multivariate bases up to arbitrary degrees and specified PCE coefficients.